# On the use of the Gram matrix for multivariate functional principal components analysis

### v1

@@ Appendix A Derivation of the equalities
$\left(L_{X}(f),u\right)_{\mathbf{M}}=\langle\!\langle f,L^{\star}_{X}(u)\rangle\!\rangle_{\Gamma},\quad\text{for all}\quad f\in\mathcal{H},\quad u\in\mathbb{R}^{N}.$ (46)
@@
So, the equality (46) is proved and we conclude that $L^{\star}_{X}$ is the adjoint operator of $L_{X}$.
Next, we derive the total inertia of the cloud $\mathcal{C}_{\!P}$ using the distance $d$. Recall that
$\operatorname{Var}\{X^{(p)}(t_{p})\}=\sum_{n=1}^{N}\pi_{n}\{X_{n}^{(p)}(t_{p})\}^{2}-\{\mu^{(p)}(t_{p})\}^{2}\quad\text{where}\quad\mu^{(p)}(t_{p})=\sum_{n=1}^{N}\pi_{n}X_{n}^{(p)}(t_{p}),\quad t_{p}\in\mathcal{T}_{p}.$ (47)
@@
The equalities in Equation (21) are shown.
@@
We now derive the total inertia of the cloud $\mathcal{C}_{\!P}$ using the distance $d_{\Gamma}$. We have that
$\sum_{n=1}^{N}\pi_{n}d_{\Gamma}^{2}(\mathrm{M}_{n},\mathrm{G}_{\!\mu})$ $=\sum_{n=1}^{N}\pi_{n}\langle\!\langle X_{n}-\mu,X_{n}-\mu\rangle\!\rangle_{\Gamma}$ (48)
$=\sum_{n=1}^{N}\pi_{n}\left|\!\left|\!\left|X_{n}\right|\!\right|\!\right|_{\Gamma}^{2}-\left|\!\left|\!\left|\mu\right|\!\right|\!\right|_{\Gamma}^{2}$ (49)
$=\sum_{p=1}^{P}\left(\sum_{n=1}^{N}\pi_{n}\left|\!\left|X_{n}^{(p)}\right|\!\right|_{\Gamma}^{2}-\left|\!\left|\mu^{(p)}\right|\!\right|_{\Gamma}^{2}\right)$ (50)
$=\sum_{p=1}^{P}\sum_{q=1}^{P}\int_{\mathcal{T}_{p}}\int_{\mathcal{T}_{q}}C_{p,q}(t_{p},s_{q})C_{p,q}(t_{p},s_{q}){\rm d}s_{q}{\rm d}t_{p}$ (51)
$=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p},$ (52)
$\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}d_{\Gamma}^{2}(\mathrm{M}_{i},\mathrm{M}_{j})$ $=\sum_{i=1}^{N}\pi_{i}\left|\!\left|\!\left|X_{i}\right|\!\right|\!\right|_{\Gamma}^{2}-\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\langle\!\langle X_{i},X_{j}\rangle\!\rangle_{\Gamma}$ (53)
$=\sum_{i=1}^{N}\pi_{i}\left|\!\left|\!\left|X_{i}\right|\!\right|\!\right|_{\Gamma}^{2}-\left|\!\left|\!\left|\mu\right|\!\right|\!\right|_{\Gamma}^{2}$ (54)
$=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}.$ (55)
The equalities in Equation (23) are shown.
@@
Finally, we derive the inertia of the cloud $\mathcal{C}_{\!N}$ using the distance $\mathsf{d}$. We have that
$\sum_{n=1}^{N}\pi_{n}\mathsf{d}^{2}(\mathsf{M}_{n},\mathsf{G}_{\!\mu})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle$ (56)
$=\sum_{p=1}^{P}\sum_{q=1}^{P}\int_{\mathcal{T}_{p}}\int_{\mathcal{T}_{q}}C_{p,q}(t_{p},s_{q})C_{p,q}(t_{p},s_{q}){\rm d}s_{q}{\rm d}t_{p}$ (57)
$=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p},$ (58)
$\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\mathsf{d}^{2}(\mathsf{M}_{i},\mathsf{M}_{j})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{n=1}^{N}\pi_{i}\pi_{j}\pi_{n}\langle\!\langle X_{n}-\mu,X_{i}-X_{j}\rangle\!\rangle\langle\!\langle X_{n}-\mu,X_{i}-X_{j}\rangle\!\rangle$ (59)
$=2\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle$ (60)
$=2\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}.$ (61)
The equalities in Equation (24) are shown.
@@ Appendix B Derivation of the eigencomponents
Using the Hilbert-Schmidt theorem, there exists a complete orthonormal basis of eigenvectors $\{v_{k}\}_{1\leq k\leq N}$ of the inner-product matrix $\mathbf{M}$ such that
$\mathbf{M}v_{k}=l_{k}v_{k}.$ (62)
Let $X=\left(X_{1}-\mu,\dots,X_{N}-\mu\right)^{\top}$ and denote $\widetilde{X}=\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}X$, the matrix of weighted observations. Recall that, in the case of $P$-dimensional process, the realisations of the process $X_{n},\leavevmode\nobreak\ n=1,\cdots,N$ and $\mu$ are vectors of functions of length $P$, and thus $X$ (and $\widetilde{X}$) is a matrix of functions of size $N\times P$. By left multiplying Equation (62) by $\widetilde{X}^{\top}$, we obtain
$\widetilde{X}^{\top}\mathbf{M}v_{k}=l_{k}\widetilde{X}^{\top}v_{k}.$ (63)
Expanding Equation (63), for each component $p=1,\dots,P$, we have,
$\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{i}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle=l_{k}\mkern-5.0mu\sum_{n=1}^{N}\mkern-4.0mu\sqrt{\pi_{n}}[v_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}.$ (64)
Here and in the following, we note $[a]_{p}$ the $p$th entry of the vector $a$. Starting from the left side of Equation (64), we get
$[\widetilde{X}^{\top}\mathbf{M}v_{k}]_{p}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{i}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle$ (65)
$=\sum_{q=1}^{P}\int_{\mathcal{T}_{q}}\sum_{i=1}^{N}\pi_{i}\left\{X_{i}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}\left\{X_{i}^{(q)}(s_{q})-\mu^{(q)}(s_{q})\right\}$ (66)
$\qquad\qquad\sum_{j=1}^{N}\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{j}^{(q)}(s_{q})-\mu^{(q)}(s_{q})\right\}{\rm d}s_{q}$ (67)
$=\sum_{q=1}^{P}\int_{\mathcal{T}_{q}}C_{p,q}(\cdot,s_{q})\sum_{j=1}^{N}\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{j}^{(q)}(s_{q})-\mu^{(q)}(s_{q})\right\}{\rm d}s_{q}$ (68)
$=\sum_{j=1}^{N}\langle\!\langle C_{p\cdot}(\cdot,\cdot),\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{j}-\mu\right\}\rangle\!\rangle$ (69)
$=\Gamma\left(\sum_{j=1}^{N}\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{j}-\mu\right\}\right)^{\mkern-9.0mu(p)}\mkern-18.0mu(\cdot),$ (70)
and, starting from the right side of Equation (64),
$[l_{k}\widetilde{X}^{\top}v_{k}]_{p}=l_{k}\sum_{n=1}^{N}\sqrt{\pi_{n}}[v_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}.$ (71)
From Equation (65) and Equation (71), we obtain
$\Gamma\left(\sum_{j=1}^{N}\sqrt{\pi_{j}}[v_{k}]_{j}\left\{X_{j}-\mu\right\}\right)^{\mkern-9.0mu(p)}\mkern-18.0mu(\cdot)=l_{k}\sum_{n=1}^{N}\sqrt{\pi_{n}}[v_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\},\quad\text{for all}\leavevmode\nobreak\ p=1,\dots,P.$ (72)
@@
$\lambda_{k}=l_{k}\quad\text{and}\quad\phi_{k}^{(p)}(\cdot)=\sum_{n=1}^{N}\sqrt{\pi_{n}}[v_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\},\quad k\geq 1.$ (73)
@@
$\left|\!\left|\!\left|\phi_{k}\right|\!\right|\!\right|^{2}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\sqrt{\pi_{i}\pi_{j}}[v_{k}]_{i}[v_{k}]_{j}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle=\sum_{i=1}^{N}[v_{k}]_{i}\sum_{j=1}^{N}\mathbf{M}_{ij}[v_{k}]_{j}$
$=\sum_{i=1}^{N}[v_{k}]_{i}l_{k}[v_{k}]_{i}=l_{k}\left|\!\left|v_{k}\right|\!\right|^{2}=l_{k}.$

### v2

@@ A.1 Proof of Lemma 1
$\left(L_{X}(f),u\right)_{\mathbf{M}}=\langle\!\langle f,L^{\star}_{X}(u)\rangle\!\rangle_{\Gamma},\quad\text{for all}\quad f\in\mathcal{H},\quad u\in\mathbb{R}^{N}.$ (A.1)
Proof.
@@
So, the equality (A.1) is proved and we conclude that $L^{\star}_{X}$ is the adjoint operator of $L_{X}$. ∎
A.2 Derivation of the inertia of the clouds
Recall that
$\operatorname{Var}\{X^{(p)}(t_{p})\}=\sum_{n=1}^{N}\pi_{n}\{X_{n}^{(p)}(t_{p})\}^{2}-\{\mu^{(p)}(t_{p})\}^{2}\quad\text{where}\quad\mu^{(p)}(t_{p})=\sum_{n=1}^{N}\pi_{n}X_{n}^{(p)}(t_{p}),\quad t_{p}\in\mathcal{T}_{p}.$ (A.2)
We first show that the total inertia of the cloud of individuals $\mathcal{C}_{\!P}$ using the distance $d$ is given by equalities in (24):
$\sum_{n=1}^{N}\pi_{n}d^{2}(\mathrm{M}_{n},\mathrm{G}_{\!\mu})=\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}d^{2}(\mathrm{M}_{i},\mathrm{M}_{j})=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\operatorname{Var}{X^{(p)}(t_{p})}{\rm d}t_{p}.$
Proof.
@@ A.2 Derivation of the inertia of the clouds
The equalities in Equation (24) are shown. ∎
@@
Next, we derive the total inertia of the cloud of individuals $\mathcal{C}_{\!P}$ using the distance $d_{\Gamma}$, given by equalities in (26):
$\sum_{n=1}^{N}\pi_{n}d_{\Gamma}^{2}(\mathrm{M}_{n},\mathrm{G}_{\!\mu})=\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}d_{\Gamma}^{2}(\mathrm{M}_{i},\mathrm{M}_{j})=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}.$
Proof.
$\sum_{n=1}^{N}\pi_{n}d_{\Gamma}^{2}(\mathrm{M}_{n},\mathrm{G}_{\!\mu})$ $=\sum_{n=1}^{N}\pi_{n}\langle\!\langle X_{n}-\mu,X_{n}-\mu\rangle\!\rangle_{\Gamma}$ (A.3)
$=\sum_{n=1}^{N}\pi_{n}\left|\!\left|\!\left|X_{n}\right|\!\right|\!\right|_{\Gamma}^{2}-\left|\!\left|\!\left|\mu\right|\!\right|\!\right|_{\Gamma}^{2}$ (A.4)
$=\sum_{p=1}^{P}\left(\sum_{n=1}^{N}\pi_{n}\left|\!\left|X_{n}^{(p)}\right|\!\right|_{\Gamma}^{2}-\left|\!\left|\mu^{(p)}\right|\!\right|_{\Gamma}^{2}\right)$ (A.5)
$=\sum_{p=1}^{P}\sum_{q=1}^{P}\int_{\mathcal{T}_{p}}\int_{\mathcal{T}_{q}}C_{p,q}(t_{p},s_{q})C_{p,q}(t_{p},s_{q}){\rm d}s_{q}{\rm d}t_{p}$ (A.6)
$=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p},$ (A.7)
$\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}d_{\Gamma}^{2}(\mathrm{M}_{i},\mathrm{M}_{j})$ $=\sum_{i=1}^{N}\pi_{i}\left|\!\left|\!\left|X_{i}\right|\!\right|\!\right|_{\Gamma}^{2}-\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\langle\!\langle X_{i},X_{j}\rangle\!\rangle_{\Gamma}$ (A.8)
$=\sum_{i=1}^{N}\pi_{i}\left|\!\left|\!\left|X_{i}\right|\!\right|\!\right|_{\Gamma}^{2}-\left|\!\left|\!\left|\mu\right|\!\right|\!\right|_{\Gamma}^{2}$ (A.9)
$=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}.$ (A.10)
The equalities in Equation (26) are shown. ∎
@@
Finally, we derive the inertia of the cloud of features $\mathcal{C}_{\!N}$ using the distance $\mathsf{d}$, given by equalities in (27):
$\sum_{n=1}^{N}\pi_{n}\mathsf{d}^{2}(\mathsf{M}_{n},\mathsf{G}_{\!\mu})=\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\mathsf{d}^{2}(\mathsf{M}_{i},\mathsf{M}_{j})=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}.$
Proof.
$\sum_{n=1}^{N}\pi_{n}\mathsf{d}^{2}(\mathsf{M}_{n},\mathsf{G}_{\!\mu})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle$ (A.11)
$=\sum_{p=1}^{P}\sum_{q=1}^{P}\int_{\mathcal{T}_{p}}\int_{\mathcal{T}_{q}}C_{p,q}(t_{p},s_{q})C_{p,q}(t_{p},s_{q}){\rm d}s_{q}{\rm d}t_{p}$ (A.12)
$=\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p},$ (A.13)
$\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\mathsf{d}^{2}(\mathsf{M}_{i},\mathsf{M}_{j})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{n=1}^{N}\pi_{i}\pi_{j}\pi_{n}\langle\!\langle X_{n}-\mu,X_{i}-X_{j}\rangle\!\rangle\langle\!\langle X_{n}-\mu,X_{i}-X_{j}\rangle\!\rangle$ (A.14)
$=2\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\pi_{j}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle$ (A.15)
$=2\sum_{p=1}^{P}\int_{\mathcal{T}_{p}}\left|\!\left|\!\left|C_{p\cdot}(t_{p},\cdot)\right|\!\right|\!\right|^{2}{\rm d}t_{p}.$ (A.16)
The equalities in Equation (27) are shown. ∎
@@ Appendix B Derivation of the eigencomponents
B.1 General framework
In this section, we calculate the relationships between the eigenelements of the covariance operator $\Gamma$ and the ones of the Gram matrix $\mathbf{M}$ of a functional dataset. We then prove the equalities (36), (37) and (38).
Using the Hilbert-Schmidt theorem, there exists a complete orthonormal basis of eigenvectors $\{\boldsymbol{u}_{k}\}_{1\leq k\leq N}$ of the inner-product matrix $\mathbf{M}$ such that
$\mathbf{M}\boldsymbol{u}_{k}=l_{k}\boldsymbol{u}_{k}.$ (A.17)
Let $X=\left(X_{1}-\mu,\dots,X_{N}-\mu\right)^{\top}$ and denote $\widetilde{X}=\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}X$, the matrix of weighted observations. Recall that, in the case of $P$-dimensional process, the realisations of the process $X_{n},~{}n=1,\cdots,N$ and $\mu$ are vectors of functions of length $P$, and thus $X$ (and $\widetilde{X}$) is a matrix of functions of size $N\times P$. By left multiplying Equation (A.17) by $\widetilde{X}^{\top}$, we obtain
$\widetilde{X}^{\top}\mathbf{M}\boldsymbol{u}_{k}=l_{k}\widetilde{X}^{\top}\boldsymbol{u}_{k}.$ (A.18)
Expanding Equation (A.18), for each component $p=1,\dots,P$, we have,
$\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\mkern-4.0mu\left\{X_{i}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle=l_{k}\mkern-5.0mu\sum_{n=1}^{N}\mkern-4.0mu\sqrt{\pi_{n}}[\boldsymbol{u}_{k}]_{n}\mkern-4.0mu\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}.$ (A.19)
Here and in the following, we note $[a]_{p}$ the $p$th entry of the vector $a$. Starting from the left side of Equation (A.19), we get
$[\widetilde{X}^{\top}\mathbf{M}\boldsymbol{u}_{k}]_{p}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\pi_{i}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\left\{X_{i}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle$ (A.20)
$=\sum_{q=1}^{P}\int_{\mathcal{T}_{q}}\sum_{i=1}^{N}\pi_{i}\left\{X_{i}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}\left\{X_{i}^{(q)}(s_{q})-\mu^{(q)}(s_{q})\right\}$ (A.21)
×$\qquad\qquad\times\sum_{j=1}^{N}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\left\{X_{j}^{(q)}(s_{q})-\mu^{(q)}(s_{q})\right\}{\rm d}s_{q}$ (A.22)
$=\sum_{q=1}^{P}\int_{\mathcal{T}_{q}}C_{p,q}(\cdot,s_{q})\sum_{j=1}^{N}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\left\{X_{j}^{(q)}(s_{q})-\mu^{(q)}(s_{q})\right\}{\rm d}s_{q}$ (A.23)
$=\sum_{j=1}^{N}\langle\!\langle C_{p\cdot}(\cdot,\cdot),\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\left\{X_{j}-\mu\right\}\rangle\!\rangle$ (A.24)
$=\Gamma\left(\sum_{j=1}^{N}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\left\{X_{j}-\mu\right\}\right)^{\mkern-9.0mu(p)}\mkern-18.0mu(\cdot),$ (A.25)
and, starting from the right side of Equation (A.19),
$[l_{k}\widetilde{X}^{\top}\boldsymbol{u}_{k}]_{p}=l_{k}\sum_{n=1}^{N}\sqrt{\pi_{n}}[\boldsymbol{u}_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\}.$ (A.26)
From Equation (A.20) and Equation (A.26), we obtain
$\Gamma\left(\sum_{j=1}^{N}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\left\{X_{j}-\mu\right\}\right)^{\mkern-9.0mu(p)}\mkern-18.0mu(\cdot)=l_{k}\sum_{n=1}^{N}\sqrt{\pi_{n}}[\boldsymbol{u}_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\},\quad\text{for all}~{}p=1,\dots,P.$ (A.27)
@@ B.1 General framework
$\lambda_{k}=l_{k}\quad\text{and}\quad\phi_{k}^{(p)}(\cdot)=\sum_{n=1}^{N}\sqrt{\pi_{n}}[\boldsymbol{u}_{k}]_{n}\left\{X_{n}^{(p)}(\cdot)-\mu^{(p)}(\cdot)\right\},\quad k\geq 1.$ (A.28)
@@
$\left|\!\left|\!\left|\phi_{k}\right|\!\right|\!\right|^{2}$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\sqrt{\pi_{i}\pi_{j}}[\boldsymbol{u}_{k}]_{i}[\boldsymbol{u}_{k}]_{j}\langle\!\langle X_{i}-\mu,X_{j}-\mu\rangle\!\rangle=\sum_{i=1}^{N}[\boldsymbol{u}_{k}]_{i}\sum_{j=1}^{N}\mathbf{M}_{ij}[\boldsymbol{u}_{k}]_{j}$
$=\sum_{i=1}^{N}[\boldsymbol{u}_{k}]_{i}l_{k}[\boldsymbol{u}_{k}]_{i}=l_{k}\left(\!\left(\boldsymbol{u}_{k}\right)\!\right)^{2}=l_{k}.$